OpenStudy (anonymous):

Find the limit for: \[\huge \lim_{x \rightarrow \infty} \frac{x + x^2}{1 − 3x^2} \] It is not where the vertical asymptote is: \[ \frac{1}{\sqrt{3}} \] And it is not 1/3 or ∞ (a medal will be given to first & best with steps and answer)

OpenStudy (experimentx):

should be -1/3

OpenStudy (anonymous):

O_O How did you see that so fast? After the first derivative this is what I have written \[\frac{1+2x}{0+6x} = \frac{1}{3}\]

OpenStudy (anonymous):

Oh crap... Never mind I see it now... -6x >_<

OpenStudy (experimentx):

divide all by x^2 take limit you will get 1/inf + 1 ----------- 1/inf - 3

OpenStudy (experimentx):

whic is -1/3